OpenStudy (anonymous):

log(x+1)=5; solve for x

OpenStudy (mimi_x3):

Write it as; \(log_e(x+1)=5\) => \((x+1)= e^{5}\)

OpenStudy (mimi_x3):

Are you able to do it now?

OpenStudy (lgbasallote):

some prefer 10 instead of e though...especially in common logarithms

OpenStudy (anonymous):

how do i type it in a calculator?

OpenStudy (helder_edwin):

@Mimi_x3 why did you assume the base of log is e?

OpenStudy (lgbasallote):

i think she's had too much of calculus =))

OpenStudy (mimi_x3):

Because it's easier to see for the RHS because the other side must be \(e\). That was what I learnt lol. @lgbasallote: it wont make sense if its \(10\)

OpenStudy (lgbasallote):

common logarithms..

OpenStudy (lgbasallote):

\[log a \implies \log_{10} a\] \[\ln a \implies \log_e a\]

OpenStudy (mimi_x3):

What is that? Care to refresh my mind. Haven't touched loagarithms in a while..

OpenStudy (mimi_x3):

It's the same thing lol

OpenStudy (mimi_x3):

log_10(a) = log_e

OpenStudy (lgbasallote):

common logarithms are logarithms with no base. it is impicit that the base is 10

OpenStudy (lgbasallote):

yes @Mimi_x3 like i said "prefer" i didnt say it was wrong

OpenStudy (mimi_x3):

why is it called "common log"

OpenStudy (mimi_x3):

wait..if you did log_10 instead; it wont make sense..

OpenStudy (lgbasallote):

http://en.wikipedia.org/wiki/Common_logarithm

OpenStudy (lgbasallote):

everything will make sense after you read that

OpenStudy (mimi_x3):

lol like i have the time to :P

OpenStudy (lgbasallote):

well you're here arent you

OpenStudy (mimi_x3):

"log10(x) should be lg (x) and loge(x) should be ln (x)."- it's the same lol